OpenStudy (anonymous):

ABCD is a parallelogram. if AB=3x+2 and CD=4x-3 find the length of AB 20 19 17 15

OpenStudy (anonymous):

Opposite sides of a parallelogram are congruent. Thus AB = CD.

OpenStudy (anonymous):

I don't understand.

OpenStudy (anonymous):

Let's see. Since opposite sides of a parallelogram are congruent and AB and CD are opposite sides, then AB = CD. Substitution gives 3x+2=4x-3 3x-4x=-3-2 -x=-5 x=5. Thus AB=3x+2=3(5)+2=17.

OpenStudy (anonymous):

Can I use that formula on any similar question?

OpenStudy (anonymous):

As long as opposite sides are given, yes. You can also set up equality using opposite angles.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

your welcome! :D